OpenStudy (anonymous):

If a 12 foot long is cut into 3/4 foot long, how many pieces will there be altogether?

OpenStudy (anonymous):

Dividing by 3/4 is the same as multiplying by 4/3.

OpenStudy (anonymous):

do you understand what they said

OpenStudy (anonymous):

my bad i mean 12/1 * 4/3 = 36/3 = 12

OpenStudy (anonymous):

i am understand

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

12 times 4/3 = 48/3 = 16.